A sealed cubical container 27.0 cm on a side contains three times Avogadro's number of molecules at a temperature of 21.0
Answer #1

We can use the ideal gas equation PV = nRT for this.
P = nRT/V
n = 3 moles, V = 0.27^3 = 0.019683 m^3, T = 294.15 K, R = 8.314472 J/(mole-K)

P = nRT/V
P = 372763.5938 Pa
F = PA =372763.5938*0.27^2 = 27174.465 N
Answer #2

number of moles = 3


PV = nRT


P*0.27^3 = 3*8.314*(21+273)


P = 372552.35 Pa


F = P*A = 372552.35*6*0.27^2 = 162.954 kN
Answer #3

no. of moles=3

T=273+21=294K

volume=19.683l

PV=nRT

P=3.67atm